The reaction 2A

B is second order with a rate constant of 51.0/M

·

min at 24

°

C.


(a) Starting with [A]0 = 9.30 ×10−3M, how long will it take for [A]t = 2.80 ×10−3M?
  min

(b) Calculate the half-life of the reaction.
min
Answer #1

for second order, a) t-takt | 5 0 x 1 | - | - 2.8x103 9.3x153 9 40 - 6- 5 xt t=4.8.9 min Ź - + =Rt. 1 -1 = kty c=0/ 3 - - you